1% solution of other compound is isotonic with 5% sucrose (sugar) solution. Then molecular wt. of compound will be

Grade:12th pass

1 Answers

Jayant Kishore
8 years ago
Dear Yogiraj,
5 % sucrose solution means 5g of sucrose in 100ml Solution ie 50g sucrose in 1 litre solution.
Molality of sucrose solution = 50/Mol Wt. = 50/342
Osmotic pressure of this solution= CRT = (50/342 ) RT
Let Mol Wt of other compound be M
Molality of unknown solution = 10/M
Osmotic Pressure of unknown solution = (10/M) RT
Assuming temperature to be same .
Since , isotonic solutions have same osmotic pressure .
=> (10/M)RT=(50/342)RT
=> M =342/5 = 68.4
So,  Mol Wt of other compound is 68.4 amu.